# Connectedness of level sets for non-degenerate integrable systems that extend complexity one torus actions

### v2

@@ Abstract.
In this paper we study the connectedness of the level sets of integrable systems that have only non-degenerate singular points and extend complexity one $T$-spaces with proper moment maps. Our main result states that if there are no singular points with a hyperbolic block and connected $T$-stabilizer, then each level set is connected. Moreover, we prove that the above condition is necessary if either some reduced space is simply connected or the moment map for the integrable system is generic in a natural sense.
@@
Keywords: Completely integrable Hamiltonian systems, non-degenerate singular points, complexity one actions, connected level sets.
@@ 1. Introduction
An important step in studying Hamiltonian group actions is understanding the level sets of the moment map. In this direction, connectedness of the level sets has been established in the following cases:
@@
In both cases, the connectedness of the level sets of the moment map is key to understanding the moment map image and plays an important role in classification.
In this paper we study the connectedness of the level sets of integrable systems that are extensions of complexity one torus actions, and have only non-degenerate singular points. Complexity one torus actions are the simplest after the toric case and have been partially classified in [KT1, KT2, KT3]. Non-degenerate singular points of integrable systems are generic, and can be thought of as the symplectic analog of critical points for Morse-Bott functions; they have attracted much interest, both for classical and quantum integrable systems (see [BF, SV] and references therein). In particular, the integrable systems that we study include the semi-toric systems of [VN], and their higher dimensional generalization in [Wa].
In our main result, Theorem 1.6 we give sufficient conditions for level sets of the moment map to be connected, generalizing the results in [VN, Wa]. Morse theory plays a central role in the theorems (1) and (2) above, and our theorem both uses and generalizes the following simple Morse-theoretic result: Let $N$ be a closed, connected 2-dimensional manifold and let $\overline{g}\colon N\to{\mathbb{R}}$ be a Morse function. If $\overline{g}$ has no singular point of index one, then each level set is connected (see Lemma 7.3). We also explore the extent to which the hypotheses in our main result are necessary (see Theorem 1.10).
@@
Here ${\mathfrak{I}}(w)$ denotes the imaginary part of $w\in{\mathbb{C}}$ and ${\mathbb{C}}^{\dim W}$ has the standard symplectic structure. If all the blocks are elliptic, we say that $p$ has purely elliptic type.
@@
The notion of non-degeneracy can be viewed as a “symplectic Morse-Bott” condition. In particular, in dimension two, a point is non-degenerate in the sense of Definition 1.2 if and only if it is non-degenerate in the sense of Morse theory. Moreover, it is hyperbolic if and only if it has index one and elliptic otherwise.
Next we introduce complexity one torus actions. Let $T$ be an $(n-1)$-dimensional torus with Lie algebra ${\mathfrak{t}}$.
@@
Let $(M,\omega,\Phi)$ be a complexity one $T$-space. Given $\alpha\in{\mathfrak{t}}^{*}$, the reduced space at $\alpha$ is the quotient $\Phi^{-1}(\alpha)/T$; if $\alpha=0$, we write ${M/\!/T}$ for the reduced space. A point $p\in M$ is short if $[p]$ is an isolated point in the reduced space $\Phi^{-1}(\Phi(p))/T$ and is tall otherwise. By [KT1, Proposition 6.1], if $\Phi$ is proper each reduced space $\Phi^{-1}(\alpha)/T$ containing more than one point is topologically a closed, connected oriented 2-dimensional manifold; moreover, any two non-empty reduced spaces have the same fundamental group [KT1, Lemma 5.7 and Corollary 9.7].
@@
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g)\right)$ be an integrable system so that $(M,\omega,\Phi)$ is a complexity one $T$-space with a proper moment map. Assume that each tall singular point in $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f\right)$ is non-degenerate. If $M$ has no non-degenerate singular points with a hyperbolic block and connected $T$-stabilizer, then the level set $f^{-1}(\alpha,c)$ is connected for each $(\alpha,c)\in{\mathfrak{t}}^{*}\times{\mathbb{R}}$; moreover, the reduced space $\Phi^{-1}(\alpha)/T$ is simply connected.
@@
Theorem 1.6 implies connectedness of the level sets for the higher dimensional analogs of semi-toric systems that Wacheux studied in [Wa]: Integrable systems $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ so that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space with a proper moment map all of whose singular points are non-degenerate without hyperbolic blocks. In particular, our main result generalizes [Wa, Theorem 4.2.4] to arbitrary dimensions.
@@
Theorem 1.6 is comparable to [PRV, Theorems 1 – 3] in that all these results assume that the singular points of an integrable system are non-degenerate, disallow (certain) singular points with a hyperbolic block, and prove that the level sets are connected. However, the results in [PRV] do not require the first component of the moment map to generate a Hamiltonian circle action, but only hold in dimension four and have additional hypotheses on the moment map.
Unfortunately, Theorem 1.6 is not quite sharp. If a reduced space has positive genus, then $M$ must have non-degenerate singular points with a hyperbolic block and connected $T$-stabilizer; see Remark 1.14. However, as we show below, it is still possible for all the level sets of $f$ to be connected in this case.
@@
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g)\right)$ be an integrable system so that $(M,\omega,\Phi)$ is a complexity one $T$-space with a proper moment map. Assume that each tall singular point in $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f\right)$ is non-degenerate. Moreover, assume that one of the following holds:
  • The reduced space $\Phi^{-1}(\beta)/T$ is simply connected for some $\beta\in\Phi(M).$
  • Each level set of $f$ contains at most one orbit of non-degenerate singular points with a hyperbolic block.
Then the level set $f^{-1}(\alpha,c)$ is connected for all $(\alpha,c)\in{\mathfrak{t}}^{*}\times{\mathbb{R}}$ if and only $M$ has no non-degenerate singular points with a hyperbolic block and connected $T$-stabilizer.
@@
If in Theorems 1.6 and 1.10 we additionally assume that $M$ is compact, then by a theorem of Hui Li [Li], the fundamental group of $M$ is equal to that of any reduced space. Hence, in this case, the phrase “the reduced space $\Phi^{-1}(\alpha)/T$” can be changed to “$M$”.
The fundamental idea behind the proofs of Theorems 1.6 and 1.10 is quite simple: Since $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f=(\Phi,g)\right)$ is an integrable system, the function $g$ is $T$-invariant, and hence induces a function $\overline{g}\colon\Phi^{-1}(\alpha)/T\to{\mathbb{R}}$. By the key technical result in our paper, Proposition 7.1, each non-trivial component of the reduced space $\Phi^{-1}(\alpha)/T$ can be given the structure of a smooth oriented surface so that $\overline{g}\colon\Phi^{-1}(\alpha)/T\to{\mathbb{R}}$ is a Morse function. Moreover, the critical points of index $1$ correspond exactly to orbits of non-degenerate singular points with a hyperbolic black and connected $T$-stabilizer. Since $\Phi$ is proper, the reduced space $\Phi^{-1}(\alpha)/T$ is compact and connected by Theorem 5.5. Therefore, the proofs reduce to standard results in Morse theory; see Lemmas 7.3 and 7.4.
@@
The second bullet point in Theorem 1.10 can be thought of as a genericity condition because generic Morse functions have at most one critical point (of index 1) in each level set.
@@
Let $(M,\omega,{\mathfrak{t}}^{*}\times{\mathbb{R}},f)$ be an integrable system satisfying the assumptions in the first two sentences of Theorem 1.6, but assume that the reduced space $\Phi^{-1}(\alpha)/T$ has positive genus $k$, and hence Euler characteristic $2-2k$. In this case, by calculating the Euler characteristic, we see that $\Phi^{-1}(\alpha)/T$ has at least $2k$ orbits of non-degenerate singular points with a hyperbolic block and connected $T$-stabilizer because because $\overline{g}$ has at least one minimum and maximum. Note that if $\Phi^{-1}(\alpha)/T$ has more than $2k$ such orbits, then the same argument shows that $\overline{g}$ must have additional local minima or maxima; hence the fibers of $f$ are not connected by Lemma 7.2.
@@
In Section 2, we review some facts about tall complexity one local models and then briefly explore a few additional properties. Next, in Section 3 we prove that we can endow the symplectic quotient $\Phi^{-1}(\alpha)/T$ with a smooth structure so that $\overline{g}$ is a Morse function – as long as most singular points satisfy a certain technical condition that guarantees that we can do this locally. Most of the rest of the paper is dedicated to proving that non-degenerate singular points satisfy this condition. For example, in Section 4 we prove that tall points that have purely elliptic type satisfy it, and in Section 5 we show that, if $M$ is four-dimensional, focus-focus points satisfy it. At this point, we are able to demonstrate our technique by using it to give a new proof that that level sets of a semi-toric system are connected. In Section 6, we prove that all the remaining types of non-degenerate singular points also satisfy this technical condition. Therefore, we can complete the proof of our theorems in Section 7.
@@ 2. Tall local models
To begin, we review some facts about tall complexity one local models. In particular, we introduce (tall) complexity one local models, construct the “defining polynomial” for tall local models, and use it to describe the sympletic quotient ${Y/\!/T}$. Moreover, we provide a generating set for $T$-invariant polynomials on these models.
@@
where $\Phi_{H}(z)=\frac{1}{2}\sum_{i}\eta_{i}|z_{i}|^{2}$ is the homogeneous moment map for the $H$-action on ${\mathbb{C}}^{h+1}$; see Remark 4.2. By the Marle-Guillemin-Sternberg local normal form theorem, there is a $T$-equivariant symplectomorphism from an invariant neighborhood of $\mathfrak{O}$ to an open subset of $Y$ taking $p$ to $[1,0,0]$. We call $(Y,\omega_{Y},\Phi_{Y})$ the local model for $p$. We say that $Y$ is short if $\Phi_{Y}^{-1}(0)$ consists of a single $T$-orbit and tall otherwise. By Lemma 2.7 below, the level set $\Phi^{-1}_{Y}(0)$ contains exactly one $T$-orbit exactly if $[p]\in\Phi^{-1}(\Phi(p))$ is an isolated point. Hence, $p$ is short exactly if its local model is short.
As shown in [KT1], tall local models have an important invariant, called the defining polynomial.
@@
Let $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a local model associated to an injective homomorphism $\rho:H\to(S^{1})^{h+1}$. Then there exists $\xi=(\xi_{0},\ldots,\xi_{h})\in{\mathbb{Z}}^{h+1}$ such that the homomorphism $(S^{1})^{h+1}\to S^{1}$ given by $\lambda\mapsto\prod\limits_{j=0}^{h}\lambda_{j}^{\xi_{j}}$ induces a short exact sequence
@@
$0\to{\mathbb{R}}^{*}\to\left({\mathbb{R}}^{h+1}\right)^{*}\stackrel{{\scriptstyle\rho^{*}}}{{\to}}{\mathfrak{h}}^{*}\to 0$
@@
Fix a tall local model [^1] Here, and throughout the paper, we only consider complexity one local models. $Y$. By Lemma 2.1, there exists a unique $\xi=(\xi_{0},\ldots,\xi_{h})\in{\mathbb{Z}}_{\geq 0}^{h+1}$ inducing the short exact sequence (2.2). The map $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$ given by $P(z)=\prod_{j=0}^{h}z_{j}^{\xi_{j}}$ is called the defining polynomial of $Y$. By a slight abuse of notation, we use the same name for the map $P\colon Y\to{\mathbb{C}}$ given by $[t,\alpha,z]\mapsto\prod_{j=0}^{h}z_{j}^{\xi_{j}}$. The sum $N:=\sum_{j=0}^{h}\xi_{j}$ is called the degree of $P$.
@@
Let ${\mathbb{Z}}_{2}=\{\pm 1\}$ act on ${\mathbb{C}}$ by $\alpha\cdot z=\alpha z$. The associated local model $S^{1}\times_{{\mathbb{Z}}_{2}}{\mathbb{R}}^{*}\times{\mathbb{C}}$ is tall. The defining polynomial $P\colon{\mathbb{C}}\to{\mathbb{C}}$ is given by $P(z)=z^{2}$ and has degree $2$.
@@
Let $S^{1}$ act effectively on ${\mathbb{C}}^{2}$ by $\alpha\cdot(z_{1},z_{2})=(\alpha z_{1},\alpha^{-1}z_{2})$ with moment map $\Phi(z_{1},z_{2})=\frac{1}{2}(|z_{1}|^{2}-|z_{2}|^{2})$. This is a tall local model. The defining polynomial $P\colon{\mathbb{C}}^{2}\to{\mathbb{C}}$ is given by $P(z_{1},z_{2})=z_{1}z_{2}$ and also has degree $2$.
@@
Following [KT1], we say that an orbit is exceptional if every nearby orbit in the same moment level set has strictly smaller stabilizer. Arguing as in the proof of Lemma 2.1, since $Y$ is tall the moment map $\Phi_{Y}$ is not proper. The first claim follows at once by [KT1, Corollary 6.3]. By [KT1, Lemmas 5.15 and 5.17], the orbit of $[1,0,0]$ is the only possible exceptional orbit in ${Y/\!/T}$. The second claim now follows at once by [KT1, Corollary 7.2]. ∎
@@
induces a diffeomorphism between ${\mathbb{C}}^{k+1}/\!/K$ and ${Y/\!/T}$. In particular, if $P(z)=z_{0}$ then $\overline{P}\colon{Y/\!/T}\to{\mathbb{C}}$ is a diffeomorphism with inverse $z\mapsto[1,0,(z,0)].$
@@
On the reduced space ${Y/\!/T}$ there is a simple relation between the magnitudes of the defining polynomial and of $z$.
@@
Let $Y:=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a tall local model with moment map $\Phi_{Y}$. Let $N$ be the degree of the defining polynomial $P(z)=\prod_{j}z_{j}^{\xi_{j}}$. Then for all $[t,\alpha,z]\in{Y/\!/T}$,
@@
Let $\eta_{0},\dots,\eta_{h}\in\mathfrak{h}^{*}$ be the weights for the $H$-action on ${\mathbb{C}}^{h+1}$. The moment map $\Phi_{Y}$ sends $[t,\alpha,z]\in Y$ to $\alpha+\Phi_{H}(z)\in{\mathfrak{t}}^{*}$, where $\Phi_{H}(z)=\sum_{j=0}^{h}\frac{1}{2}|z_{j}|^{2}\eta_{j}$. Thus it suffices to check that $\lvert z\rvert^{2}=\big{(}\prod_{j}\xi_{j}^{\xi_{j}}\big{)}^{-\frac{1}{N}}\big{\lvert}P(z)\big{\rvert}^{\frac{2}{N}}$ for all $z\in\Phi_{H}^{-1}(0)$. Fix $z\in\Phi_{H}^{-1}(0)$. Since the local model $Y$ is tall and $\Phi_{H}(z)=\sum_{j=0}^{h}\frac{1}{2}|z_{j}|^{2}\eta_{j}=0$, by Remark 2.3 there exists $c\geq 0$ such that $|z_{j}|^{2}=c\xi_{j}$ for all $j$. Therefore,
@@
Let $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a tall local model with moment map $\Phi_{Y}\colon Y\to{\mathfrak{t}}^{*}$. The algebra of $T$-invariant polynomials on $Y$ is generated by the real and imaginary parts of the defining polynomial $P\colon Y\to{\mathbb{C}}$, the map $[t,\alpha,z]\mapsto|z|^{2}$, and the components of $\Phi_{Y}$.

### v3

# Connectedness of fibers beyond semitoric systems I: the non-degenerate case

@@ Abstract.
In this paper we study the connectedness of the fibers of integrable systems that have only non-degenerate singular points and extend complexity one $T$-spaces with proper moment maps. Our main result states that if there are no tall singular points with a hyperbolic block and connected $T$-stabilizer, then each fiber is connected. Moreover, we prove that the above condition is necessary if either some reduced space is simply connected or the moment map for the integrable system is generic in a natural sense.
@@
Keywords: Completely integrable Hamiltonian systems, non-degenerate singular points, complexity one actions, connected fibers.
@@ 1. Introduction
An important step in studying Hamiltonian group actions is understanding the fibers of the moment map. In this direction, connectedness of the fibers has been established in the following cases:
@@
In both cases, the connectedness of the fibers of the moment map is key to understanding the moment map image and plays an important role in classification.
In this paper we study the connectedness of the fibers of integrable systems that are extensions of complexity one torus actions, and have only non-degenerate singular points. Complexity one torus actions are the simplest after the toric case and have been partially classified in [KT1, KT2, KT3]. Non-degenerate singular points of integrable systems are generic, and can be thought of as the symplectic analog of critical points for Morse-Bott functions; they have attracted much interest, both for classical and quantum integrable systems (see [BF, SV] and references therein). In particular, the integrable systems that we study include the semi-toric systems of [VN], and their higher dimensional generalization in [Wa].
In our main result, Theorem 1.6 we give sufficient conditions for fibers of the moment map to be connected, generalizing the results in [VN, Wa]. Morse theory plays a central role in the theorems (1) and (2) above, and our theorem both uses and generalizes the following simple Morse-theoretic result: Let $N$ be a closed, connected 2-dimensional manifold and let $\overline{g}\colon N\to{\mathbb{R}}$ be a Morse function. If $\overline{g}$ has no singular point of index one, then each fiber is connected (see Lemma 7.3). We also explore the extent to which the hypotheses in our main result are necessary (see Theorem 1.10).
@@
Here ${\mathfrak{I}}(w)$ denotes the imaginary part of $w\in{\mathbb{C}}$ and ${\mathbb{C}}^{\dim W}$ has the standard symplectic structure. Moreover, $p$ has purely elliptic type if $p$ is non-degenerate and all the blocks are elliptic.
@@
The notion of non-degeneracy can be viewed as a “symplectic Morse-Bott” condition. In particular, in dimension two, a singular point is non-degenerate in the sense of Definition 1.2 if and only if it is non-degenerate in the sense of Morse theory. Moreover, it is hyperbolic if and only if it has index one and elliptic otherwise.
Next we introduce complexity one torus actions. Let $T$ be an (n1)(n-1)-dimensional torus with Lie algebra $\mathfrak{t}$}.
@@
Let $(M,\omega,\Phi)$ be a complexity one $T$-space. Given $\beta\in{\mathfrak{t}}^{*}$, the reduced space at $\beta$ is the quotient $\Phi^{-1}(\beta)/T$; if $\beta=0$, we write M//T{M/\!/T} for the reduced space. A point $p\in M$ is short if $[p]$ is an isolated point in the reduced space $\Phi^{-1}(\Phi(p))/T$ and is tall otherwise. By [KT1, Proposition 6.1], if $\Phi$ is proper each reduced space $\Phi^{-1}(\beta)/T$ containing more than one point is topologically a closed, connected oriented 2-dimensional manifold; moreover, any two non-empty reduced spaces have the same fundamental group [KT1, Lemma 5.7 and Corollary 9.7].
@@
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g)\right)$ be an integrable system such that $(M,\omega,\Phi)$ is a complexity one $T$-space with a proper moment map. Assume that each tall singular point in $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f\right)$ is non-degenerate and that no such point has a hyperbolic block and connected $T$-stabilizer. Then the fiber $f^{-1}(\beta,c)$ is connected for each $(\beta,c)\in{\mathfrak{t}}^{*}\times{\mathbb{R}}$; moreover, the reduced space $\Phi^{-1}(\beta)/T$ is simply connected.
@@
Theorem 1.6 implies connectedness of the fibers for the higher dimensional analogs of semi-toric systems that Wacheux studied in [Wa]: Integrable systems $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ such that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space with a proper moment map all of whose singular points are non-degenerate without hyperbolic blocks. In particular, our main result generalizes [Wa, Theorem 4.2.4] to arbitrary dimensions.
@@
Theorem 1.6 is comparable to [PRV, Theorems 1 – 3] in that all these results assume that the singular points of an integrable system are non-degenerate, disallow (certain) singular points with a hyperbolic block, and prove that the fibers are connected. However, the results in [PRV] do not require the first component of the moment map to generate a Hamiltonian circle action, but only hold in dimension four and have additional hypotheses on the moment map.
Unfortunately, Theorem 1.6 is not quite sharp. If a reduced space has positive genus, then $M$ must have non-degenerate singular points with a hyperbolic block and connected $T$-stabilizer; see Remark 1.14. However, as we show below, it is still possible for all the fibers of $f$ to be connected in this case.
@@
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g)\right)$ be an integrable system such that $(M,\omega,\Phi)$ is a complexity one $T$-space with a proper moment map. Assume that each tall singular point in $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f\right)$ is non-degenerate. Moreover, assume that one of the following holds:
  • The reduced space $\Phi^{-1}(\beta_{0})/T$ is simply connected for some $\beta_{0}\in\Phi(M).$
  • Each fiber of $f$ contains at most one $T$-orbit of non-degenerate singular points with a hyperbolic block.
Then the fiber $f^{-1}(\beta,c)$ is connected for all $(\beta,c)\in{\mathfrak{t}}^{*}\times{\mathbb{R}}$ if and only $M$ has no tall singular points with a hyperbolic block and connected $T$-stabilizer.
@@
If in Theorems 1.6 and 1.10 we additionally assume that $M$ is compact, then by a theorem of Hui Li [Li], the fundamental group of $M$ is equal to that of any reduced space. Hence, in this case, the phrase “the reduced space $\Phi^{-1}(\beta)/T$” can be changed to “MM”.
The fundamental idea behind the proofs of Theorems 1.6 and 1.10 is quite simple: Since $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f=(\Phi,g)\right)$ is an integrable system, the function $g$ is $T$-invariant, and hence induces a function $\overline{g}\colon\Phi^{-1}(\beta)/T\to{\mathbb{R}}$. By the key technical result in our paper, Proposition 7.1, each non-trivial component of the reduced space $\Phi^{-1}(\beta)/T$ can be given the structure of a smooth oriented surface so that $\overline{g}\colon\Phi^{-1}(\beta)/T\to{\mathbb{R}}$ is a Morse function. Moreover, the critical points of index $1$ correspond exactly to orbits of non-degenerate singular points with a hyperbolic black and connected $T$-stabilizer. Since $\Phi$ is proper, the reduced space $\Phi^{-1}(\beta)/T$ is compact and connected by Theorem 5.5. Therefore, the proofs reduce to standard results in Morse theory; see Lemmas 7.3 and 7.4.
@@
The second bullet point in Theorem 1.10 can be thought of as a genericity condition because generic Morse functions have at most one critical point (of index 1) in each fiber.
@@
Let $(M,\omega,{\mathfrak{t}}^{*}\times{\mathbb{R}},f)$ be an integrable system satisfying the assumptions in the first two sentences of Theorem 1.6, but assume that the reduced space $\Phi^{-1}(\beta)/T$ has positive genus $k$, and hence Euler characteristic 22k2-2k. In this case, by calculating the Euler characteristic, we see that $\Phi^{-1}(\beta)/T$ has at least $2k$ orbits of non-degenerate singular points with a hyperbolic block and connected $T$-stabilizer because because $\overline{g}$ has at least one minimum and maximum. Note that if $\Phi^{-1}(\beta)/T$ has more than $2k$ such orbits, then the same argument shows that $\overline{g}$ must have additional local minima or maxima; hence the fibers of $f$ are not connected by Lemma 7.2.
@@
In Section 2, we review some facts about tall complexity one local models and then briefly explore a few additional properties. Next, in Section 3 we prove that we can endow the symplectic quotient $\Phi^{-1}(\beta)/T$ with a smooth structure so that $\overline{g}$ is a Morse function – as long as most singular points satisfy a certain technical condition that guarantees that we can do this locally. Most of the rest of the paper is dedicated to proving that non-degenerate singular points satisfy this condition. For example, in Section 4 we prove that tall points that have purely elliptic type satisfy it, and in Section 5 we show that, if $M$ is four-dimensional, focus-focus points satisfy it. At this point, we are able to demonstrate our technique by using it to give a new proof that that fibers of a semi-toric system are connected. In Section 6, we prove that all the remaining types of non-degenerate singular points also satisfy this technical condition. Therefore, we can complete the proof of our theorems in Section 7.
@@ 2. Tall local models
To begin, we review some facts about tall complexity one local models. In particular, we introduce (tall) complexity one local models, construct the “defining polynomial” for tall local models, and use it to describe the symplectic quotient Y//T{Y/\!/T}. Moreover, we provide a generating set for $T$-invariant polynomials on these models.
@@
where $\Phi_{H}(z)=\frac{1}{2}\sum_{i}\eta_{i}|z_{i}|^{2}$ is the homogeneous moment map for the $H$-action on ${\mathbb{C}}^{h+1}$; see Remark 4.2. By the Marle-Guillemin-Sternberg local normal form theorem, there is a $T$-equivariant symplectomorphism from an invariant neighborhood of $\mathfrak{O}$ to an open subset of $Y$ taking $p$ to $[1,0,0]$. We call $(Y,\omega_{Y},\Phi_{Y})$ the local model for $p$. We say that $Y$ is short if $\Phi_{Y}^{-1}(0)$ consists of a single $T$-orbit and tall otherwise. By Lemma 2.7 below, the fiber $\Phi^{-1}_{Y}(0)$ contains exactly one $T$-orbit exactly if $p]\in\Phi^{-1}(\Phi(p))/T$ is an isolated point. Hence, $p$ is short exactly if its local model is short.
As shown in [KT1], tall local models [^1] Here, and throughout the paper, all the local models that we consider are complexity one. have an important invariant, called the defining polynomial.
@@
Let $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a local model associated to an injective homomorphism $\rho\colon H\to(S^{1})^{h+1}$. Then there exists $\xi=(\xi_{0},\ldots,\xi_{h})\in{\mathbb{Z}}^{h+1}$ such that the homomorphism $(S^{1})^{h+1}\to S^{1}$ given by $\lambda\mapsto\prod\limits_{j=0}^{h}\lambda_{j}^{\xi_{j}}$ induces a short exact sequence
@@
$0\to{\mathbb{R}}^{*}\to\left({\mathbb{R}}^{h+1}\right)^{*}\stackrel{{\scriptstyle\rho^{*}}}{{\to}}{\mathfrak{h}}^{*}\to 0.$
@@
Fix a tall local model $Y$. By Lemma 2.1, there exists a unique $\xi=(\xi_{0},\ldots,\xi_{h})\in{\mathbb{Z}}_{\geq 0}^{h+1}$ inducing the short exact sequence (2.2). The map $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$ given by $P(z)=\prod_{j=0}^{h}z_{j}^{\xi_{j}}$ is called the defining polynomial of $Y$. By a slight abuse of notation, we use the same name for the map $P\colon Y\to{\mathbb{C}}$ given by $t,\alpha,z]\mapsto\prod_{j=0}^{h}z_{j}^{\xi_{j}}$. The sum $N:=\sum_{j=0}^{h}\xi_{j}$ is called the degree of $P$.
@@
Let ${\mathbb{Z}}_{2}=\{\pm 1\}$ act on $\mathbb{C}$} by $\lambda\cdot z=\lambda z$. The associated local model $S^{1}\times_{{\mathbb{Z}}_{2}}{\mathbb{R}}^{*}\times{\mathbb{C}}$ is tall. The defining polynomial $P\colon{\mathbb{C}}\to{\mathbb{C}}$ is given by $P(z)=z^{2}$ and has degree $2$.
@@
Let $S^{1}$ act effectively on ${\mathbb{C}}^{2}$ by $\lambda\cdot(z_{1},z_{2})=(\lambda z_{1},\lambda^{-1}z_{2})$ with moment map $\Phi(z_{1},z_{2})=\frac{1}{2}(|z_{1}|^{2}-|z_{2}|^{2})$. This is a tall local model. The defining polynomial $P\colon{\mathbb{C}}^{2}\to{\mathbb{C}}$ is given by $P(z_{1},z_{2})=z_{1}z_{2}$ and also has degree $2$.
@@
Following [KT1], we say that an orbit is exceptional if every nearby orbit in the same moment fiber has strictly smaller stabilizer. Arguing as in the proof of Lemma 2.1, since $Y$ is tall the moment map $\Phi_{Y}$ is not proper. The first claim follows at once by [KT1, Corollary 6.3]. By [KT1, Lemmas 5.15 and 5.17], the orbit of $[1,0,0]$ is the only possible exceptional orbit in Y//T{Y/\!/T}. The second claim now follows at once by [KT1, Corollary 7.2]. ∎
@@
induces a diffeomorphism between ${\mathbb{C}}^{k+1}/\!/K$ and Y//T{Y/\!/T}. In particular, if $P(z)=z_{0}$ then $\overline{P}\colon{Y/\!/T}\to{\mathbb{C}}$ is a diffeomorphism with inverse $z\mapsto[1,0,(z,0)].$
@@
On the reduced space Y//T{Y/\!/T} there is a simple relation between the magnitudes of the defining polynomial and of $z$.
@@
Let $Y:=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a tall local model with moment map $\Phi_{Y}$. Let $N$ be the degree of the defining polynomial $P(z)=\prod_{j}z_{j}^{\xi_{j}}$. Then for all $t,\alpha,z]\in{Y/\!/T}$,
@@
Let $\eta_{0},\dots,\eta_{h}\in\mathfrak{h}^{*}$ be the weights for the $H$-action on ${\mathbb{C}}^{h+1}$. The moment map $\Phi_{Y}$ sends $t,\alpha,z]\in Y$ to $\alpha+\Phi_{H}(z)\in{\mathfrak{t}}^{*}$, where $\Phi_{H}(z)=\sum_{j=0}^{h}\frac{1}{2}|z_{j}|^{2}\eta_{j}$. Thus it suffices to check that $\lvert z\rvert^{2}=\big(\prod_{j}\xi_{j}^{\xi_{j}}\big)^{-\frac{1}{N}}\big\lvert P(z)\big\rvert^{\frac{2}{N}}$ for all $z\in\Phi_{H}^{-1}(0)$. Fix $z\in\Phi_{H}^{-1}(0)$. Since the local model $Y$ is tall and $\Phi_{H}(z)=\sum_{j=0}^{h}\frac{1}{2}|z_{j}|^{2}\eta_{j}=0$, by Remark 2.3 there exists $c\geq 0$ such that $|z_{j}|^{2}=c\xi_{j}$ for all $j$. Therefore,
@@
Let $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a tall local model with moment map $\Phi_{Y}\colon Y\to{\mathfrak{t}}^{*}$. The algebra of $T$-invariant polynomials on $Y$ is generated by the real and imaginary parts of the defining polynomial $P\colon Y\to{\mathbb{C}}$, the map $t,\alpha,z]\mapsto|z|^{2}$, and the components of $\Phi_{Y}$.